# Enhanced Detection of Tiny Objects in Aerial Images

### v1

@@ 4.5 Attention Mechanism
Augmented Backbone $_{50} APAP$ Recall Precision
@@
Default 0.286 0.142 0.262 0.547
1st Design 0.289 0.143 0.271 0.527
2nd Design 0.283 0.142 0.264 0.503
3rd Design 0.274 0.131 0.258 0.577
4th Design 0.336 0.177 0.301 0.625
5th Design ✓ 0.344 0.180 0.301 0.687
6th Design 0.342 0.178 0.315 0.593
Dataset: VisDrone (150 epochs)
5th Design ✓ 0.360 0.216 0.347 0.467
6th Design 0.358 0.214 0.352 0.453
Table 6: Backbone designs
[Uncaptioned image] $(640\times 640\times 3)$ $(640\times 640\times 3)$
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] $(320\times 320\times 64\times w)$ $(320\times 320\times 128\times w)$
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] $(160\times 160\times 128\times w)$ $(160\times 160\times 256\times w)$
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] $(80\times 80\times 256\times w)$ $(80\times 80\times 512\times w)$
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] $(40\times 40\times 512\times w)$ $(40\times 40\times 1024\times w)$
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] $(20\times 20\times 1024\times w)$ $(20\times 20\times 2048\times w)$
4.6 Final training and ablation studies
The final MoonNet-based model was trained from scratch with optimized input image resolutions and augmentations, showing clear improvements over the default YOLOv8n. In Table 7, Default refers to the original backbone, Res to optimized input image resolution, and Aug to optimized augmentations. The number of channels from the model’s backbone in the final training is not identical to the numbers presented in table 5. This is because all the final trainings in the table 7 were conducted using YOLOv8n model scale thus actual number of channels were (32, 64, 128, 256, 512) due to the channel multiplier (0.25). For the fair comparison, all the backbones in table 7 used same number of channels.
@@ 4.6 Final training and ablation studies
Method 1 0.397 0.236 0.356 0.638
Method 2 0.439 0.268 0.387 0.680
Method 3 0.441 0.268 0.375 0.766
Method 4 0.504 0.299 0.442 0.781
Image resolution adjustment (Method1 $\to$ Method2)
Varying the input image resolution yields +10.58% gain in $AP_{50}$ and +13.56% gain in $AP$ showing improved optimization helps recover missed tiny objects.
Proper data augmentation (Method2 $\to$ Method3)
Adding geometric transformations such as rotation and flip steadies generalisation: $AP_{50}$ rises slightly to 0.441, while $AP$ stays the same.
Effectiveness of MoonNet (Method3 $\to$ Method4)
Replacing the vanilla CSP with MoonNet achieves the highest scores: 0.504 $AP_{50}$ and 0.299 $AP$. This yields +14.29% gain and +11.57% gain over the tuned backbone, and +26.91% gain in $AP_{50}$ over the baseline.
@@
YOLC(k=1) [12] HRNet + MoonNet(Ours) 0.329 0.550 0.336
In table 8, the MoonNet’s attention module arrangement was integrated into the YOLC model and achieved state-of-the-art performance when compared with ClusDet[21] and DMNet[22], both trained on cluster-aware cropped images. Ours obtained +15.87% gain compared to ClusDet, +11.9%gain compared to DMNet and +8.58% gain against the original YOLC.
Table 9: Visual comparison
Method 1 Method 2 Method 3
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Method 4 YOLC YOLC + MoonNet
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
@@ 5 Conclusion
This research improves tiny object detection by combining resolution adjustment, augmentation, and the MoonNet backbone. Compared to default YOLOv8, the final model (Model 4) achieved gains of +26.96% $AP_{50}$, +26.69% $AP$, +24.16% recall, and +22.41% precision. In addition, the MoonNet integrated backbone achieved state-of-the-art performance among the cluster-aware detectors confirming attention-augmented backbone enhances detection of small targets in aerial imagery.

### v2

@@ 4.5 Attention & gating mechanism
Backbone sigmoid 1+tanh $_{50} APAP$ Recall Precision
@@
Default no attention 0.486 0.299 0.427 0.690
1st Design 0.491 0.301 0.437 0.685
0.471 0.287 0.417 0.700
2nd Design 0.489 0.299 0.433 0.702
0.467 0.279 0.415 0.702
3rd Design 0.503 0.313 0.444 0.697
0.490 0.300 0.430 0.702
4th Design 0.483 0.294 0.441 0.681
0.473 0.289 0.427 0.680
5th Design 0.483 0.293 0.425 0.706
0.472 0.286 0.424 0.682
6th Design 0.485 0.297 0.434 0.685
0.471 0.282 0.422 0.677
Dataset: VisDrone (50 epochs)
1st Design 0.294 0.170 0.303 0.414
2nd Design 0.296 0.172 0.304 0.412
3rd Design 0.293 0.169 0.303 0.409
4.6 Final training and integration with YOLC
The final MoonNet-based model was trained from scratch with optimized input image resolutions and augmentations, showing clear improvements over the default YOLOv8n. In Table 5, Default refers to the original backbone, Res to optimized input image resolution, and Aug to optimized augmentations. The number of channels from the model’s backbone in the final training is not identical to the numbers presented in Table 4. This is because all the final trainings in the Table 5 were conducted using YOLOv8n model scale thus actual number of channels were (16, 32, 64, 128, 256) due to the channel multiplier (0.25). In Table 6, all parameters from preprocess to end-to-end are measured in milliseconds.
@@ 4.6 Final training and integration with YOLC
Method 1 0.491 0.308 0.390 0.685
Method 2 0.596 0.411 0.501 0.763
Method 3 0.664 0.485 0.565 0.843
Method 4 0.667 0.486 0.566 0.849
Table 6: GFLOPs and latency comparison
Method GFLOPs Preprocess Inference Postprocess End-to-End
Method 1 8.3 0.1 1.9 2.3 4.3
Method 2 8.3 0.2 2.2 2.6 5.0
Method 3 8.3 0.2 2.2 2.7 5.1
Method 4 8.4 0.2 2.6 2.4 5.2
The MoonNet’s (SE Block–CBAM) orientation is then adapted to the YOLC framework to check the MoonNet’s adaptability and enhancement potential. MoonNet (sigmoid) and MoonNet (1+tanh) both achieved improved results compare to the original YOLC in Table 7. The MoonNet (1+tanh) achieved +6.13% gain in $AP$ and +3.77% gain in $AP_{50}$ compare to the MoonNet (sigmoid).
@@
YOLC(k=1)(Ours) HRNet + MoonNet(sigmoid) 0.310 0.530 0.311
YOLC(k=1)(Ours) HRNet + MoonNet(1+tanh) 0.329 0.550 0.336
Table 8: Gating mechanism comparison
Framework sigmoid 1+tanh $_{50} APAP$
YOLO-MoonNet (DOTA) 0.503 0.313
YOLO-MoonNet (DOTA) 0.490 0.300
YOLC-MoonNet (VisDrone) 0.310 0.530
YOLC-MoonNet (VisDrone) 0.329 0.550
Method 1 Method 2 Method 3 Method 4 YOLC YOLC+MoonNet
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Fig. 3: Visual comparison of detection performance on tiny objects across two different aerial scenes.
4.7 Ablation Study
Image resolution adjustment (Method1 $\to$ Method2)
Varying the input image resolution yields +21.38% gain in $AP_{50}$ and +33.44% gain in $AP$ showing improved optimization helps recover missed tiny objects.
Proper data augmentation (Method2 $\to$ Method3)
Instead of adding excessive geometric transformations such as rotation and color space adjustment, mild use of augmentations leads to +11.41% gain in $AP_{50}$ and +18% gain in $AP$.
Effectiveness of MoonNet (Method3 $\to$ Method4)
Compared to the previous two steps, replacement of the backbone from vanilla CSP to MoonNet yields lower increment. +0.45% gain in $AP_{50}$ and +0.21% gain in $AP$, but obtained the highest scores: 0.667 $AP_{50}$ and 0.486 $AP$.
Computational burden (GFLOPs and latency)
For the training using the final configuration, only 0.1 increase in GFLOPs and 0.9 ms increase were measured from the base score in Table 6 which highlights MoonNet’s high deployability with small costs.
Hybrid use of attention modules (SE Block & CBAM)
Our third design MoonNet which consists of multiple attention modules demonstrated significant enhancement in detection performance. However, the power of multi-use of attention modules is still questionable. In Table 4, design 4, 5 and 6 also used SE Block and CBAM together but failed to achieve higher metrics than the designs with single-type attention module. MoonNet’s SE Block orientation in the early layer and CBAM in the later stage conveys that the channel emphasis from SE Block followed by spatial focus from CBAM can be particularly effective for combined use of multiple attention modules. When using more than one attention module within a single backbone, careful consideration in module arrangement is required.
Adaptability of MoonNet (YOLC integration)
In Table 7, the MoonNet’s attention module arrangement was integrated into the YOLC model with two different gating functions and achieved state-of-the-art performance when compared with ClusDet[21] and DMNet[7], both trained on cluster-aware cropped images. Ours(sigmoid) obtained +16.10% gain in $AP$ compared to ClusDet, +8.77% gain compared to DMNet and +2.31% gain against the original YOLC. Ours(1+tanh) obtained +23.22% gain in $AP$ compared to ClusDet, +15.44% gain compared to DMNet and +8.58% gain against the original YOLC.
Alternative gating mechanism ($1+tanh()$)
An attention module’s gating mechanism is framework-dependent and this is evident through Table 8’s result. In YOLC framework (HRNet), identity safe gating using tanh might be a better option. Our ablation also revealed that the identity safe gating is unnecessary for the YOLO (CSPDarknet) framework. Therefore, careful consideration on choosing gate function is essential.
4.8 Visual predictions
Visual predictions were made in Figure 3. As the method evolves from 1 to 4, the model begins to capture more small vehicles, and the method 4 model correctly performs separate detection of small and large vehicles. YOLC+MoonNet model also demonstrates better detection performance.
@@ 5 Conclusion
This research improves tiny object detection by combining resolution adjustment, augmentation, and the MoonNet backbone with tanh gating function. Compared to the base YOLOv8n-obb model, the final model (Method 4) achieved gains of +35.85% $AP_{50}$, +57.79% $AP$, +45.13% recall, and +23.94% precision. Moreover, the MoonNet-integrated backbone achieved state-of-the-art performance among the cluster-aware detectors confirming an attention-augmented backbone enhances detection of small targets in aerial imagery. However, the multiple use of attention modules cannot guarantee consistent enhancement. Its orientation and gating function need to be tuned for the appropriate implementation.